Sub 24
Ques. 15

A 20 m ladder and a 15 m ladder were leaned against a building. The bottom of the longer ladder was 7 m farther from the building than the bottom of the shorter ladder, but both ladders are the same distance up the building. Find the distance.
1
Expert's answer
2012-05-31T08:24:22-0400
x^2+y^2 = 73
4x+4y = 44, x +y = 11, x = 11- y

(11 - y)^2 +y^2 = 73,
y^2 -11y +24 = 0, y1 = 3, y2 = 8
x1 = 8, x2 = 3.
So the sides are 3 ft and
8 ft